# Mamo: a Mathematical Modeling Benchmark with Solvers

### v2

@@ Appendix B Cross Review
In the optimization section, particularly with easy LP part, out of 688 entries, 12 were found to be completely invalid (not meeting the criteria), 26 had incorrect answers, and 6 had misleading descriptions. All invalid entries were filtered out, and the remaining issues were corrected manually. Among them, 8 of them were corrected and others (36 questions) were filtered out. In complex LP part, for the original 211 questions, 40 were found formulated by wrong mathematical models. All of them were corrected.
@@ B.1 Cross-Review Process
Furthermore, adhering to the scaling law evident in the dataset’s distribution, we selected 50 questions from the ODE parts in our dataset for a deeper analysis. Four independent reviewers (see their qualification in Section C) were tasked with assessing 50 questions each   [^10] The hourly wage pay is 100 RMB/h, the total amount spent on the participant compensation is 2400 RMB, approximately 331 dollars. The responses could either be numeric or labeled as “error” if a question was deemed problematic. This section outlines the metrics used to evaluate the reviewers’ responses against pre-defined correct answers, which also include the “error” label for any flawed questions. We use the same metric as described in Subsection 6.1.2. Here are our results:
@@ B.1.1 Results
Appendix C Annotator Qualifications
@@ C.1 Qualification of Reviewers
Appendix D Metric
@@
$\text{Correctness}=\begin{cases}1&\text{if }\left|\frac{O-A}{A}\right|\leq 1\times 10^{-4}\text{~{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} (1)} or }|O-A|<1\times min\{10^{-n},10^{-2}\}\text{~{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} (2)},}\\ 0&\text{otherwise.}\end{cases}$
@@
(2) The only difference lies in the last few digits of the required significant figure of the answer. By selecting different thresholds $\xi\times min\{10^{-n},10^{-2}\}$, we can adjust the tolerance for the discrepancy, ensuring that $|O-A|<\xi\times min\{10^{-n},10^{-2}\}$. We have explored the relationship between various thresholds and the accuracy of GPT-4o.
@@
Figure 2: Accuracy of different thresholds in different categories
@@
Appendix E Error Analysis
@@
E.1 Deviation from Instruction
@@
E.2 Syntax Error
In ODE parts, the syntax errors is obviously the syntax errors in the reponsed python code. In Optimization part, the .lp file require more strict syntax than the universal optimization formulation. The Figure 4 and 4 show the typical syntax errors in .lp format.
@@ E.3 Modeling Error
Appendix F Scope of Mamo
Table 7: The taxonomy of mathematical modeling. In the Mamo benchmark, we only include the categories where sophisticated solvers are available.
@@ Appendix F Scope of Mamo
Appendix G Examples of Different ${\widehat{M}}$
@@
Appendix H Examples of Syntax Errors
@@
However, if we replace the constraint $c1$ with x <= 20 - y, the code will be reported as an error, even if the two constraints are equivalent. To specifically test the modelling ability of the LLMs, we apply code modifier to fix the syntax errors.
Appendix I Few-shot Experiment
@@
Appendix J Testing Process
We use 3-shot learning in testing; see examples in Appendix O. The examples of testing process in ODE and optimization is shown in Figure 8 and Figure 9.
@@
Appendix K Question Formulation
Figure  10 shows the example of synthesis process in ODE.
@@
Appendix L Word Cloud
See the word cloud of the combined data in Figure 11
@@
Appendix M Evaluation script
@@
  • We testing the following models by calling API: All the closed source models, DeepSeek-V2 and DeepSeek-coder , with temperature=0.8
@@
Appendix O Prompts
@@
[Instruction] Assume you are a virtual assistant with expertise in optimization, specifically in creating .lp files for linear programming problems. Your task is to translate given natural language problems into optimization models, formatted as .lp files.When you receive a question, it might include mathematical expressions in LaTeX format. Your job is to interpret these expressions accurately and model the problem in .lp format. Your response must adhere to the following guidelines: - The optimization model must be written in .lp format, adhering to the conventions and syntax appropriate for linear programming problems. - The model should be designed so that solving it yields the optimal value, which directly answers the question posed. - Your response should be an entire .lp file, ready to be processed by a linear programming solver. Ensure that the file contains no comments or extraneous content beyond the model itself. - Handle LaTeX expressions with care to ensure that the mathematical aspects of the problem are accurately represented in the .lp model. - If the solution needs to be rounded to an integer, make use of the ’General’ integer constraint in the .lp file to specify integer variables, please do not use ’General’ if there are not requirement for the integer variables. Here comes the examples: [Example_1] (the input) A manufacturing company produces two types of products: $X$ and $Y$. The production cost for each unit of product $X$ is $50$, while the cost for each unit of product $Y$ is $10$. There are constraints in the production process, such that twice the number of units produced from product $X$, plus the number of units from product $Y$, cannot exceed 200 due to resource limitations. In addition, to meet market demand, four times the number of units produced from product $X$, plus the number of units from product $Y$, must be at least 50. Considering these constraints and given that both products can only be produced in whole numbers due to their physical nature, what is the minimum total cost needed for producing these items while satisfying all conditions? Provide your answer rounded to the nearest dollar. Your response: Minimize obj: 50 x + 10 y Subject To c1: 2 x + y <= 200 c2: 4 x + y >= 50 Bounds x >= 0 y >= 0 Generals x y End […] Please craft the .lp file according to these instructions, focusing on delivering a model that is directly solvable to obtain the answer. And Please follow the syntax like examples to write the .lp file. Here comes the question: [question] Generate the contents of an .lp file for this problem, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality . Ensure there is a space between variables and their coefficients. Your response:
@@
[Instruction] Assume you are a virtual assistant with expertise in ordinary differential equations (ODEs), particularly in formulating ODE models from natural language descriptions and solving them using Python’s `$solve\_ivp`$, `$odeint`$ from SciPy, and `$dsolve`$ from SymPy. Your primary task is to convert the given natural language problems into ODE models and then apply ODE solvers to find the solutions. When you receive a question, it will often contain mathematical expressions in LaTeX format, which you need to interpret accurately. Your response must be in the form of Python code that directly outputs the final solution to the problem upon execution. This Python code should adhere to the following criteria: 1. The output should only display the final answer of the problem. 2. Utilize a ’final-state-approach’ where the code immediately prints the solution after solving the ODE, without showing any intermediate steps. 3. Ensure the answer is rounded to the specified number of significant figures or decimal places. Use Python’s ‘round‘ function for decimal rounding. For significant figures, include and use the following function in your code: “‘ python def round_to_significant_figures(num, sig_figs): if num != 0: return round(num, -int(math.floor(math.log10(abs(num))) + (1 - sig_figs))) else: return 0 # Handles the case of num being 0 “‘ 4. Your response should be entirely in Python code, formatted to run directly without modifications. 5. Handle LaTeX expressions in the problem statement carefully to ensure accurate modeling and solution. Here comes the [examples] : […] Please process the problem according to these instructions, focusing solely on delivering the Python code that meets these requirements. And Please directly generate the code without any explaintion(except the comments in the code). the following lines are forbidden: Here’s the Python code that solves the given problem and meets the specified requirements: “‘ python “‘ This code uses the ‘solve_ivp’ function from SciPy to solve the initial value problem for the given differential equation. The ‘round_to_significant_figures’ function is included to round the final answer to the specified number of significant figures. Upon execution, the code will directly output the amount of pollutant left in the tank after 5 minutes, rounded to five significant figures. Please avoid the above sentences. Take a deep breathe before answering the question. This is a piece of cake to you. Here comes the question: [question] Your response:
@@
Appendix P Limitations
While the current benchmarking methodology necessitates LLMs to generate Python code for ODEs and .lp files for optimization problems, it opens up exciting avenues for future research. The process, while effective, may inadvertently influence the LLMs’ focus towards formalization over the conceptual modeling, potentially affecting the depth of mathematical reasoning. This presents an opportunity to explore innovative benchmark designs that can better distinguish between an LLM’s modeling prowess and its formalization abilities. Future research could aim to develop methodologies that allow LLMs to demonstrate their mathematical modeling capabilities without the intermediary step of formalization, leading to a more nuanced and comprehensive assessment of their abilities
Appendix Q Ethics Statement:

### v3

@@ Appendix B Cross Review
In the optimization section, particularly with easy LP part, out of 688 entries, 12 were found to be completely invalid (not meeting the criteria), 26 had incorrect answers, and six had misleading descriptions. All invalid entries were filtered out, and the remaining issues were corrected manually. Among them, eight of them were corrected and others (36 questions) were filtered out. In complex LP part, for the original 211 questions, 40 were found formulated by wrong mathematical models. All of them were corrected.
@@ B.1 Cross-Review Process
Furthermore, adhering to the scaling law evident in the dataset’s distribution, we selected 50 questions from the ODE parts in our dataset for a deeper analysis. Four independent reviewers (see their qualification in Section C) were tasked with assessing 50 questions each   [^5] The hourly wage is approximately 20 US dollars (equivalent in local currency), with a total participant compensation of around 473 US dollars (equivalent in local currency).. The responses could either be numeric or labeled as “error” if a question was deemed problematic. This section outlines the metrics used to evaluate the reviewers’ responses against pre-defined correct answers, which also include the “error” label for any flawed questions. We use the same metric as described in Subsection 4.1.2. Here are our results:
@@ B.1.1 Results
Appendix C Collector Qualifications
@@ C.1 Qualification of Reviewers
Appendix D Understand Code Modifier
Figure 5 shows the evaluation process involving the code modifier. The introduction of code modifiers is to maintain the fairness of the evaluation: we try to reduce conditions when the LLM outputs the code with a correct mathematical model, but the code ends with a compile error due to the limit of coding ability. We concider two factors when applying code modifier:
  1. (a)
    How often does a format error occur?
  2. (b)
    How many passes on average are needed?
The factor (a) is exactly the Table 5. As for (b), the number of passes required to fix format errors depends on the code modifier chosen. To study the factor (b), To study factor (b), we applied GPT-4-0613 as a code modifier on the chosen format error files and iteratively called it until the code was executed successfully. Table 7 displays the distribution of the number of pass the code modifier was called until the errors were fixed. The chosen files were generated by GPT-4o and DeepSeek-v2.5, representing a close-source and an open-source model, respectively.
1 2 3 4 $\geq$5 all
LP 74 35 10 6 11 136
ODE 88 3 1 1 9 112
Table 7: Number of pass to apply code modifier
On average, 1.68 passes were needed to correct format errors when using GPT-4-0613. We found that a single pass was sufficient to correct the majority of format errors. To balance efficiency and accuracy, we only applied a single pass for the code modifier throughout our experiments.
Appendix E Metric
@@
[…]
@@
(2) The only difference lies in the last few digits of the required significant figure of the answer. By selecting different thresholds $\xi\times min\{10^{-n},10^{-2}\}$, we can adjust the tolerance for the discrepancy, ensuring that $|\widehat{A}-A|<\xi\times min\{10^{-n},10^{-2}\}$. We have explored the relationship between various thresholds and the accuracy of GPT-4o.
@@
Figure 6: Accuracy of different threshold in different category
@@
Appendix F Error Analysis
@@
F.1 Deviation from Instruction
@@
F.2 Syntax Error
In ODE parts, the syntax errors is obviously the syntax errors in the reponsed python code. In Optimization part, the .lp file require more strict syntax than the universal optimization formulation. The Figure 8 and 8 show the typical syntax errors in .lp format.
@@ F.3 Modeling Error
Appendix G Scope of Mamo
@@
Table 9: The taxonomy of mathematical modeling. In Mamo benchmark, we only include the categories where sophisticated Solvers is available.
@@
Appendix H Examples of Different ${\widehat{M}}$
@@
Appendix I Examples of Syntax Errors
@@
However, if we replace the constraint $c1$ with x <= 20 - y, the code will be reported as an error, even if the two constraints are equivalent. To specifically test the modeling ability of the LLMs, we apply code modifier to fix the syntax errors.
Appendix J Example of Mathematical Model Correction
In Figure 10, we display an example of correction of the mathematical model in Complex_LP. In the optimization model of the max flow problem, variables were initially set in the wrong signs in two constraints($C_{1},C_{3}$). The corrected constraints($C_{1}^{\prime},C_{3}^{\prime}$) are also displayed.
[…]
Figure 10: The example of model correction
Appendix K Few-shot Experiment
@@
Appendix L Testing Process
We use 3-shot learning in testing; see examples in Appendix R. The examples of testing process in ODE and optimization is shown in Figure 13 and Figure 14.
@@
Appendix M Probability of Getting Correct $\widehat{A}$ from Wrong ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\widehat{M}}$
There are chances for the wrong mathematical model $\widehat{M}$ to lead to correct answer $\widehat{A}$. However, the probability is low due to the wide spread of the continuous answer space and lack of prior knowledge. We examined the distribution of all the answers in Mamo dataset. The probability of guessing an ODE answer within a 0.01 interval is about 2.9%, and for LP, it’s around 7.3%. These low probabilities highlight the difficulty of correctly guessing the answer without proper modeling.
Appendix N Question Formulation
Figure  15 shows the example of synthesis process in ODE. The process of synthesizing data starts with the creation of typical mathematical constructs, into which random parameters are introduced to give them specificity. This is followed by the computation of answers to establish a ground truth. Subsequently, GPT-4 is employed to craft real-life scenarios based on these predefined models, effectively translating abstract mathematical concepts into tangible, context-rich problems.
@@
Appendix O Data Description and Word Cloud
We present a comparative table showcasing the number of scenarios across various categories, from manufacturing to natural science, in Table 10. We also make the word cloud of the combined data in Figure 16
Category Manufacturing Environment Business Daily Life Natural Science Others
LP 75 114 265 258 44 77
ODE 85 74 54 22 109 2
Table 10: Comparison of ODE and LP dataset of different categories
@@
Appendix P Evaluation script
The evaluation script is as following:
@@
  • We tested the following models by calling API: All the closed source models except o1-preview, DeepSeek-V2.5, Llama-3.1 series, Qwen-2.5 series with temperature=0.8. o1-preview was called under default temperature since the change of temperature is not allowed.
@@
Appendix R Prompts
The following figures show the prompts we use in the benchmarking. Figure 17,18 display the prompt we use in the evaluation process, while Figure 19,20 list the prompts we use for calling code modifiers. As shown in Figure 19,20, the code modifier is prompted to adjust the code based solely on the code and error information during execution, without any information about the problem itself. This ensures that the code modifier does not alter the underlying mathematical model based on problem details.
@@
[Instruction] Assume you are a virtual assistant with expertise in optimization, specifically in creating .lp files for linear programming problems. Your task is to translate given natural language problems into optimization models, formatted as .lp files.When you receive a question, it might include mathematical expressions in LaTeX format. Your job is to interpret these expressions accurately and model the problem in .lp format. Your response must adhere to the following guidelines: - The optimization model must be written in .lp format, adhering to the conventions and syntax appropriate for linear programming problems. - The model should be designed so that solving it yields the optimal value, which directly answers the question posed. - Your response should be an entire .lp file, ready to be processed by a linear programming solver. Ensure that the file contains no comments or extraneous content beyond the model itself. - Handle LaTeX expressions with care to ensure that the mathematical aspects of the problem are accurately represented in the .lp model. - If the solution needs to be rounded to an integer, make use of the ’General’ integer constraint in the .lp file to specify integer variables, please do not use ’General’ if there are not requirement for the integer variables. Here comes the examples: [Example_1:] (the input) A manufacturing company produces two types of products: $X$ and $Y$. The production cost for each unit of product $X$ is $50$, while the cost for each unit of product $Y$ is $10$. There are constraints in the production process, such that twice the number of units produced from product $X$, plus the number of units from product $Y$, cannot exceed 200 due to resource limitations. In addition, to meet market demand, four times the number of units produced from product $X$, plus the number of units from product $Y$, must be at least 50. Considering these constraints and given that both products can only be produced in whole numbers due to their physical nature, what is the minimum total cost needed for producing these items while satisfying all conditions? Provide your answer rounded to the nearest dollar. Your response: Minimize obj: 50 x + 10 y Subject To c1: 2 x + y <= 200 c2: 4 x + y >= 50 Bounds x >= 0 y >= 0 Generals x y End […] Please craft the .lp file according to these instructions, focusing on delivering a model that is directly solvable to obtain the answer. And Please follow the syntax like examples to write the .lp file. Here comes the question: [question] Generate the contents of an .lp file for this problem, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality . Ensure there is a space between variables and their coefficients. Your response:
@@
[Instruction] Assume you are a virtual assistant with expertise in ordinary differential equations (ODEs), particularly in formulating ODE models from natural language descriptions and solving them using Python’s `$solve\_ivp`$, `$odeint`$ from SciPy, and `$dsolve`$ from SymPy. Your primary task is to convert the given natural language problems into ODE models and then apply ODE solvers to find the solutions. When you receive a question, it will often contain mathematical expressions in LaTeX format, which you need to interpret accurately. Your response must be in the form of Python code that directly outputs the final solution to the problem upon execution. This Python code should adhere to the following criteria: 1. The output should only display the final answer of the problem. 2. Utilize a ’final-state-approach’ where the code immediately prints the solution after solving the ODE, without showing any intermediate steps. 3. Ensure the answer is rounded to the specified number of significant figures or decimal places. Use Python’s ‘round‘ function for decimal rounding. For significant figures, include and use the following function in your code: “‘ python def round_to_significant_figures(num, sig_figs): if num != 0: return round(num, -int(math.floor(math.log10(abs(num))) + (1 - sig_figs))) else: return 0 # Handles the case of num being 0 “‘ 4. Your response should be entirely in Python code, formatted to run directly without modifications. 5. Handle LaTeX expressions in the problem statement carefully to ensure accurate modeling and solution. Here comes the examples: [Examples] […] Please process the problem according to these instructions, focusing solely on delivering the Python code that meets these requirements. And Please directly generate the code without any explaintion(except the comments in the code). the following lines are forbidden: Here’s the Python code that solves the given problem and meets the specified requirements: “‘ python “‘ This code uses the ‘solve_ivp’ function from SciPy to solve the initial value problem for the given differential equation. The ‘round_to_significant_figures’ function is included to round the final answer to the specified number of significant figures. Upon execution, the code will directly output the amount of pollutant left in the tank after 5 minutes, rounded to five significant figures. Please avoid the above sentences. Take a deep breathe before answering the question. This is a piece of cake to you. Here comes the question: [question] Your response:
@@
Appendix S Licensing Information
This content(incolding benchmark) is licensed under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) license.
Appendix T Explanation of Criteria
In this section, we explain the reasoning behind the validation of each proposition shown in Figure 3, based on the properties outlined in our benchmark dataset.
T.1 Property 1: Final-State Approach Solvability
Property 1 states that problems should be solvable directly, using a ’final-state approach’. This means that once the model is formulated, the solver should be able to compute the final answer without requiring additional transformations. We assess two propositions based on this property:
  • Proposition 1: $A=y(t_{0})$ is valid as it asks for the value of the solution function $y(t)$ at a specific time $t_{0}$, which is a final-state solution. A solver can compute this directly.
  • Proposition 1: $A=a\times y(t_{0})+b$ is invalid because it requires an additional transformation (i.e., multiplying by $a$ and adding $b$), which does not adhere to the final-state solvability approach.
T.2 Property 2: Unified and Numerical Answers
Property 2 requires the answer to be a unified and numerical value, particularly for ODE problems where the function’s value at a specific time is sought. For optimization problems, the optimal value should be clear.
  • Proposition 2: $A=\max(x+y)$ for $x,y\leq 0$ = 0.0 is valid as it yields a clear numerical result, specifically the maximum of $x+y$ within the defined constraints.
  • Proposition 2: $A=\sqrt{3}$ is invalid because while $\sqrt{3}$ represents a number, it is not expressed as a numerical answer. In practice, this could lead to ambiguities if different solvers provide decimal approximations (e.g., $1.732$) instead of the exact expression.
T.3 Property 3: Significant Figures and Precision
Property 3 mandates that each question specifies the required precision or number of significant figures for the answer.
  • Proposition 3: $A=6.66$ is valid because it presents the number to three significant figures, which is explicitly required.
  • Proposition 3: $A=6.7$ is invalid as it is given only to two significant figures, which does not meet the required precision.
T.4 Property 4: Real-World Problem Context
Finally, Property 4 introduces the need for real-world context in problem framing, without explicitly stating the mathematical model. This requires the solver to abstract the appropriate mathematical model from the problem description.
  • Proposition 4: "In a company…" aligns with this property, as the problem is framed as a real-world scenario. The mathematical model is not explicitly given, challenging the solver to formulate it based on the provided context.